OpenStudy (anonymous):

solve for x: x:5sqrtx

OpenStudy (solomonzelman):

what is exactly the equation?

OpenStudy (anonymous):

\[\frac{ x }{ 5\sqrt{x} }\]

OpenStudy (anonymous):

=24

OpenStudy (anonymous):

\[\frac{ x }{ 5\sqrt{x} }=24\]

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{x }{5\sqrt{x}}=24 }\) \(\large\color{black}{ \displaystyle 5\sqrt{x}=24 }\) \(\large\color{black}{ \displaystyle \sqrt{x}=4.8 }\) \(\large\color{black}{ \displaystyle (\sqrt{x})^2=(4.8)^2 }\) \(\large\color{black}{ \displaystyle x=? }\)

OpenStudy (anonymous):

why is \[\frac{ x }{ 5\sqrt{x}}=24 \] the same as\[5\sqrt{x}=24\]

OpenStudy (solomonzelman):

I did it incorrectly at the 3rd line

OpenStudy (solomonzelman):

to answer your question. \(\large\color{black}{ \displaystyle \frac{x}{5\sqrt{x}}=24 }\) \(\large\color{black}{ \displaystyle \frac{x\times \sqrt{x}}{5\sqrt{x}\times \sqrt{x}}=24 }\) \(\large\color{black}{ \displaystyle \frac{x\times \sqrt{x}}{5x}=24 }\) \(\large\color{black}{ \displaystyle \frac{\sqrt{x}}{5}=24 }\)

OpenStudy (solomonzelman):

then multiply both sides times 5 \(\large\color{black}{ \displaystyle \sqrt{x}=100 }\)

OpenStudy (solomonzelman):

and if \(\large\color{black}{ \displaystyle \sqrt{x}=100 }\), then \(\large\color{black}{ \displaystyle x=? }\)

OpenStudy (anonymous):

1000, so in that case 120x120

OpenStudy (anonymous):

so the answer is 14400 right ?

OpenStudy (solomonzelman):

x=10000

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sqrt{x}=100 }\) \(\large\color{black}{ \displaystyle (\sqrt{x})^2=(100 )^2 }\) \(\large\color{black}{ \displaystyle x=10{\tiny~~}000 }\)

OpenStudy (anonymous):

But 24x5=120 right. So x must be 120x120=14400 ?